OpenStudy (highschoolmom2010):

find value of x 60 57 29 33

OpenStudy (highschoolmom2010):



OpenStudy (anonymous):

you would have to use the law of sines. Or the law of cosines

OpenStudy (highschoolmom2010):

cosine I belive

OpenStudy (highschoolmom2010):

\[\cos X=\frac{ 11 }{ 20 }\]

OpenStudy (highschoolmom2010):

then this is where I get tuck

OpenStudy (anonymous):

from there you would divide both sides by cosine

OpenStudy (anonymous):

if you do it on a calculator make sure you are in degrees not radians

OpenStudy (highschoolmom2010):

divide by cosine what though

OpenStudy (anonymous):

actually i think you would arc sin the 11/20

OpenStudy (highschoolmom2010):

yea because I need to do the inverse to get the angle

OpenStudy (anonymous):

ya thats right i think

OpenStudy (highschoolmom2010):

i put in cos^-1 (11/20)=56.6329870308

OpenStudy (john_es):

It's Ok, this is the correct solution.

OpenStudy (john_es):

But you have to approximate so it will be 57.

OpenStudy (highschoolmom2010):

thanks :D @John_ES &@unwiser